# Filtering for Creativity: Adaptive Prompting for Multilingual Riddle Generation in LLMs

### v1

@@ 4 AOF Pretrained Evaluations
We evaluate riddle generation across three pretrained models using five prompting selecting riddle one example per method. This section blends automatic metrics with qualitative observations to assess metaphor, creativity, and syntactic variation. Representative riddles are provided in Appendices DG. For detailed quantitative results—token length, Self-BLEU, Distinct-2, and syntactic validity—see Appendix B.
4.1 English
GPT-4o
achieves moderate repetition (Self-BLEU: 0.413) as shown in Table 5 and high lexical diversity (Distinct-2: 0.852), balancing structural cohesion with surface novelty. Compared to LLaMA 3.1 (0.471 / 0.727) and DeepSeek R1 (0.339 / 0.845), its generation reflects a middle ground: less repetitive than LLaMA, but more structurally conventional than R1. The riddle in Row 1 of Table 6 incorporates contrastive metaphor in a coherent frame, supporting findings that figurative ambiguity paired with consistent syntax enhances interpretability Lakoff & Johnson (1980); Shutova (2013). This reflects controlled creativity, where lexical risk is constrained by structural regularity Binsted (1996).
@@ 4.1 English
displays the strongest phrasal variation among the three models, as evidenced by a higher Distinct-2 (0.727), though with a moderate increase in repetition (Self-BLEU: 0.471). In Row 2, the riddle demonstrates rhythmic structure and concept layering, leveraging symmetry and internal metaphor to guide reader inference. This aligns with cognitive accounts linking memorable riddles to rhythmic and conceptual salience Koestler (1964). The AOF method appears to help LLaMA 3.1 decouple fixed templates from overused lexical forms, allowing for recompositional variety without semantic drift Fauconnier & Turner (2002).
DeepSeek R1
exhibits the lowest repetition (Self-BLEU: 0.339) and highest token diversity (Distinct-2: 0.845), outperforming both GPT-4o and LLaMA 3.1 in novelty. The output in Row 3 exemplifies conceptual inversion and abstract framing, using oppositional imagery to support lateral interpretation. This mirrors classic riddle mechanics involving duality and semantic misdirection Koestler (1964). While diversity of this magnitude sometimes correlates with fluency degradation Zhang et al. (2021), AOF appears to regulate R1’s output sufficiently to retain syntactic legibility Xu et al. (2018), enabling expressive reformulations without loss of coherence.
4.2 Japanese
GPT-4o
GPT-4o’s performance on metrics like self-BLEU and Distinct-n under AOF ranks around average relative to standard baselines—reflecting more on metric saturation than prompting inadequacy (Yao et al. (2025), Schmidtová et al. (2024)). AOF resolves key flaws in traditional prompts, such as the egocentric “I”-imagery in chain-of-thought and overfitting in few-shot examples. These gains are not fully captured by current metrics. As shown in Table 10, GPT-4o under AOF adopts a distinct structure: a punchy first sentence followed by a more elaborate second, enhancing narrative pacing and engagement.
LLaMA 3.1
While LLaMA 3.1 shows limited improvement on automated metrics using AOF, this reflects model expressiveness more than prompt quality. AOF mitigates egocentric phrasing and rote repetition seen in baseline prompts, yielding subtler gains beyond surface metrics. For instance, Table 10 features a clever use of the homophone 「つる」 to suggest both decorative twine and the crane (鶴)—symbols deeply embedded in Japanese culture and Shinto ritual, such as しめ縄 (shimenawa) An (2023).
DeepSeek R1
DeepSeek R1’s middling scores on self-BLEU and Distinct-n under AOF say more about its stylistic tendencies than the prompt’s design Li et al. (2024). AOF corrects core issues in baseline methods—overuse of first-person voice in chain-of-thought and example mimicry in few-shot—enabling richer creativity not fully reflected in metrics. Table 10 presents a vivid example: a fish’s mouth described as a “quiet tree” where birds sing, merging the surreal and natural in a poetically disorienting twist.
4.3 Arabic
GPT-4o
GPT-4o shows moderate repetition (Self-BLEU: 0.497) as shown in Table 5 and good lexical variety (Distinct-2: 0.780) with Adaptive Originality Filtering (AOF), clearly performing better than common methods like few-shot, zero-shot, chain-of-thought, and adversarial prompts. Unlike chain-of-thought prompts, which tend to produce straightforward, predictable metaphors, AOF helps GPT-4o create riddles with imaginative and abstract images—such as something that’s present but unseen—as illustrated in (Figure 2, Row 1). This approach fits naturally with traditional Arabic riddles, known for their symbolic and reflective style Al-Khatib (1988).
LLaMA 3.1
LLaMA 3.1 strikes an effective balance between repetition (Self-BLEU: 0.374) and creativity (Distinct-2: 0.927) through AOF, addressing issues often found in chain-of-thought and adversarial prompts, which frequently yield predictable or overly vague outputs. Its riddles are relatable and culturally resonant, using clear metaphors drawn from everyday life, like ”a strong wind” that can’t enter a house, as shown in (Figure 2, Row 2). This connects directly to familiar poetic traditions in Arabic, avoiding common pitfalls like repetitive phrasing or loss of meaning Al-Jahiz (869).
DeepSeek R1
DeepSeek R1, while somewhat repetitive (Self-BLEU: 0.585), achieves notable depth in metaphorical expression (Distinct-2: 0.583) under AOF. This method effectively tackles problems seen in zero-shot, few-shot, and adversarial prompting, such as repetitive or simplistic metaphors. For example, DeepSeek R1 creatively portrays a rooftop as an eye ”fed by the city,” as seen in (Figure 2, Row 3), mixing urban imagery with striking visual symbolism. This clever blending of abstract ideas and real-world images strongly aligns with Arabic poetry, known for its layers of meaning and subtle metaphors Al-Marzouki (2012). By encouraging culturally rich riddles, AOF clearly boosts the originality and depth of DeepSeek R1’s outputs compared to simpler prompting strategies Xu et al. (2018).
4.4 French
GPT-4o
GPT-4o’s pretrained riddles are grammatically solid and easy to understand, but they often feel like lifted translations of English puzzles. For example, it offers the riddle in Row 1 of Table 17, a fluent but familiar “cloud” trope. The phrasing remains straightforward, with minimal use of inversion or enjambment that one might expect in classic French énigmes Chan (1996). Even when prompted for more creativity, GPT-4o tends to default back to elemental imagery—wind, water, shadows—rather than exploring urban or abstract concepts. Answerability is never in doubt, but the surface novelty and cultural resonance remain modest. This is reflected in a Self-BLEU of 0.413 and Distinct-2 of 0.852 as shown in Table 5, suggesting moderate repetition and relatively strong lexical variety.
DeepSeek R1
DeepSeek R1 generates concise and structurally consistent riddles, yet it leans heavily on classic “river” or “echo” formats. A typical example, shown in Row 2, feels like a direct adaptation of childhood puzzles, with little lexical or rhythmic innovation Meulemans (2005). Attempts to push deeper—such as referencing time, memory, or abstract concepts—often collapse back into familiar patterns. Even in adversarial or chain-of-thought modes, R1 seldom ventures beyond these safe metaphors. The result is always coherent but predictable, with cultural idioms and advanced metaphorical shifts under-utilized. This pattern yields a low Self-BLEU of 0.339 and a high Distinct-2 of 0.845, indicating minimal paraphrastic overlap and strong surface novelty.
LLaMA 3.1
LLaMA 3.1 shows the greatest stylistic range—some riddles stumble through literal phrasing, while others introduce intriguing wordplay. Row 3 plays with poetic misdirection (“danse” / “ris”), while Row 4 experiments with digital metaphor (“curseur”), reflecting a willingness to stretch genre boundaries. These flashes of originality are countered by examples that revert to stilted syntax or literal templates. Overall, LLaMA’s pretrained outputs blend moments of genuine creativity with occasional lapses into generic phrasing, suggesting strong potential but inconsistent performance. The variation is mirrored in its Self-BLEU score of 0.471 and Distinct-2 of 0.727, balancing moderate repetition with decent surface diversity Veale (2011).
4.5 Chinese
GPT-4o
GPT-4o’s pretrained Chinese riddles are grammatically correct and logically coherent, but they often translate English metaphors without adapting to the script-specific strategies typical of traditional 灯谜. As shown in Row 1 of Table 13, the imagery is literal and binary, missing multi-layered allusions like radical-based clues or idiomatic rhythm Chan (1996); Sun (2006). When prompted for variation, GPT-4o maintains syntactic fluency but rarely ventures into lexical innovation. With a Self-BLEU of 0.280 and Distinct-2 of 0.869 as in shown in Table 5, it balances mild repetition with surface diversity but lacks deeper cultural anchoring.
DeepSeek R1
DeepSeek R1 outputs elegant, fluent couplets with strong adherence to classical poetic symmetry, as shown in Row 2. While it maintains strong rhythm and antithesis, its metaphors remain shallow—favoring form over lexical novelty. Even when attempting radical-based clues (Row 3), the effort often feels literal rather than layered. Despite this, R1 exhibits high diversity (Distinct-2: 0.674) and moderate repetition (Self-BLEU: 0.433), suggesting a capacity for expressive phrasing that falls short of fully exploiting Chinese script-level wordplay Xu et al. (2018).
LLaMA 3.1
LLaMA 3.1 demonstrates the richest cultural range. Row 4 blends visual and semantic metaphor in a style reminiscent of folk riddles, while Row 5 shows explicit use of radical-based structure. This script-aware design reflects deeper integration with Chinese morphological conventions Li (2008). Although some outputs revert to generic “无…能…” templates or awkward logic, LLaMA’s outputs remain stylistically diverse (Distinct-2: 0.776) and moderately novel (Self-BLEU: 0.428). This suggests that the AOF prompt enables meaningful variation while retaining cultural fidelity Fauconnier & Turner (2002).
Language Pair Prompting Method Fine-Tuned GPT-4o (Self-BLEU / Distinct-2)
English–Arabic Few-Shot 0.233 / 0.826
AOF (Ours) 0.260 / 0.893
Zero-Shot 0.391 / 0.752
Chain-of-Thought 0.326 / 0.831
Adversarial 0.320 / 0.810
English–Chinese AOF (Ours) 0.163 / 0.934
Zero-Shot 0.315 / 0.831
Few-Shot 0.349 / 0.787
Chain-of-Thought 0.305 / 0.828
Adversarial 0.400 / 0.757
English–Japanese AOF (Ours) 0.177 / 0.915
Zero-Shot 0.431 / 0.752
Few-Shot 0.326 / 0.778
Chain-of-Thought 0.386 / 0.796
Adversarial 0.327 / 0.748
English–French Chain-of-Thought 0.256 / 0.892
AOF (Ours) 0.273 / 0.856
Zero-Shot 0.289 / 0.867
Few-Shot 0.323 / 0.835
Adversarial 0.359 / 0.793
Table 1: Prompting performance (Self-BLEU / Distinct-2) for the fine-tuned GPT-4o model across language pairs. One bolded cell per language pair shows the best combined performance (lowest Self-BLEU and highest Distinct-2).
@@ 4.5 Chinese
We compare GPT-4o before and after fine-tuning across five prompting strategies. Quantitative metrics—token length, Self-BLEU, and Distinct-2—are complemented by qualitative analysis of metaphorical framing, structural variation, and bilingual phrasing. Representative pairs are shown in Appendices DG.
5.1 English
[…]
5.2 Japanese
[…]
5.3 Chinese
[…]
5.4 Arabic
[…]
5.5 French
[…]
6 Fine-Tuned Riddle Comparison to Real World
We evaluate fine-tuned riddle generations across five prompting strategies by comparing them to their closest literary or folkloric counterparts. This comparison focuses on structural fidelity, metaphorical depth, and cultural alignment. Language-specific analyses are provided in Appendix A, which details how fine-tuned outputs differ from real-world riddles in phrasing, abstraction, and stylistic form.
@@ 7 Conclusion
This paper introduced Adaptive Originality Filtering (AOF), a prompting framework designed to overcome originality limits in bilingual riddle generation. Evaluated across four language pairs, five prompting paradigms, and three LLMs, AOF consistently outperformed baselines in lexical diversity (Distinct-n), semantic novelty (Self-BLEU), and cross-lingual alignment. Fine-tuned GPT-4o with AOF produced the highest-quality outputs, outperforming both its pretrained version and open-weight models. Human evaluations confirmed that AOF-generated riddles were more metaphorically rich, structurally coherent, and culturally resonant—especially in French, Japanese, Chinese, and Arabic. These findings demonstrate that filtering-based prompting, particularly when combined with fine-tuning, provides a robust path toward semantically novel and linguistically sound multilingual generation. AOF marks a practical advance in prompt design for creative, cross-lingual NLP tasks.
@@
We restrict our study to the BiRdQA dataset, covering 6,614 English and 8,751 Chinese riddles with multiple-choice answers. While genre-diverse, its figurative focus narrows generalizability to broader creative tasks (e.g., storytelling, allegory). Our multilingual evaluation spans five languages (EN–ZH–AR–JA–FR), excluding lower-resource or morphologically rich ones like Finnish or Swahili.
@@
We fix sampling parameters (e.g., temperature, token limit) across all models to ensure comparability, potentially masking prompt–parameter effects. AOF’s MiniLM-based filtering emphasizes surface-level novelty, risking semantic redundancy undetected by cosine distance—particularly in non-English outputs.
@@
Our fine-tuning on GPT-4o uses BiRdQA’s multiple-choice framing, enhancing structural fluency but biasing toward riddle types favoring clarity over ambiguity. Although stylistic gains are evident in token length and Self-BLEU, we omit extrinsic human metrics like puzzle difficulty or solver accuracy.
@@
Human evaluations cover only five languages due to cost, limiting cultural generalizability. Syntactic validity checks apply solely to AOF generations, leaving other strategies unchecked structurally. Lastly, while AOF excels in novelty-structure balance, deeper human validation (e.g., creativity ratings, solve rates) remains future work.
@@
This work evaluates riddle generation across five languages—English, Chinese, Japanese, Arabic, and French—selected for linguistic diversity and resource availability. However, our dataset and prompts reflect biases inherent to internet-derived corpora and may not fully capture the cultural or idiomatic richness of underrepresented communities. For example, certain metaphorical forms or idiomatic patterns may be overrepresented in English and underdeveloped in Arabic or Japanese, despite efforts to balance qualitative evaluation across languages.
@@
Riddles generated in this study may closely resemble publicly available riddles from folklore, educational resources, or crowdsourced riddle repositories. Although our Adaptive Originality Filtering (AOF) method actively screens for semantic overlap, we acknowledge that latent training data exposures or structural mimicry may inadvertently echo human-created content. As such, we recommend that users and downstream applications of our system avoid commercial use or publication of model outputs without further originality assessment or attribution checks.
@@
The BiRdQA dataset used for training and evaluation contains no personally identifiable information (PII). All riddles are anonymized and synthetically framed as general-knowledge metaphors. Our fine-tuning process adheres to OpenAI’s guidelines for API use and operates within approved token limits and safety thresholds. No user data or private content was used in training, generation, or evaluation.
@@
While riddle generation appears benign, the creative ambiguity inherent in figurative tasks can be exploited for misinformation or manipulation, especially in politically or culturally sensitive contexts. Models capable of abstract generation may also fabricate plausible but misleading information. We caution against deploying this system in high-stakes educational, psychological, or legal settings without adequate interpretability safeguards or human oversight.
Acknowledgements
We thank Bakr Bouhaya for his thoughtful analysis of Arabic riddles, which provided valuable insights and cultural grounding for our study.
Appendix A Appendix: Fine-Tuned AOF Riddle Comparison to Real World

### v3

@@ 6 AOF Pretrained Evaluations
Pre-trained AOF prompts improve riddle quality across all languages by promoting metaphorical novelty and structural fluency, even without fine-tuning. Cross-lingually, DeepSeek R1 consistently yields the highest RiddleScores (e.g., English: 0.400; Arabic: 0.400; Chinese: 0.453; Japanese: 0.475), suggesting strong compatibility with the AOF sampling rejection framework. These outputs combine lexical diversity with controlled syntactic rhythm (Koestler, 1964; Xu et al., 2018). For example, DeepSeek’s Arabic riddle in Figure 6, Row 3 metaphorically compares a rooftop to an “eye fed by the city,” demonstrating culturally grounded abstraction (Al-Marzouki, 2012).
DeepSeek R1
attains the highest Riddlescore in four of five languages: EN (0.400), AR (0.400), ZH (0.453), JA (0.379) - outperform GPT 4o / LLaMA 3.1 by 5-15 points (Table 4). While slightly more repetitive in Arabic (Self-BLEU: 0.585), R1 compensates with high lexical diversity—e.g., Distinct-2 scores of 0.845 in English and 0.674 in Chinese (Table 8)—and fluent metaphorical abstraction. Its Japanese riddle (Table 14, Row 3) showcasing the kind of poetic misdirection that aligns with high RiddleScore evaluations (Xu et al., 2018).
GPT‑4o
performs consistently in languages with moderate repetition (Self‑BLEU $\approx$ 0.41–0.50; Table 8), high lexical variety (Distinct‑2: 0.78–0.85), and RiddleScore values from 0.373 (FR/AR) to 0.453 (ZH) (Table 4), reflecting fluent but less figuratively ambitious riddles. Notably, in FR and ZH, GPT‑4o exhibits literal translation tendencies that limit cultural nuance (Chan, 1996; Sun, 2006).
@@
shows stylistic risk‑taking (Distinct‑2 $\approx$ 0.727–0.927; Table 8) but variable cohesion (RiddleScore: 0.330–0.378; Table 4), often blending innovative metaphors with uneven syntax or logical drift. For example, its JA riddle in Table 14 cleverly puns on the homophone tsuru (鶴/twine), linking cultural symbols via Shinto imagery (An, 2023).
Despite varied outputs, shared patterns emerge: AOF avoids template reuse, minimizes egocentric phrasing, and achieves cultural competence without tuning. These patterns, supported by Tables 4 and 8, validate the language‑agnostic nature of the metaphor‑rich generation. For complete evaluations, see Section B.
7 AOF Fine-Tuned Evaluations
Fine-tuning GPT‑4o with AOF consistently enhances riddle quality for EN, ZH, FR, and AR by improving semantic creativity, lexical variation, and cultural mastery. The increases in RiddleScore range from 33.4% (AR) to 48.3% (ZH), as shown in Table 4. Self‑BLEU reduces by 33–51% (Table 9), and Distinct‑2 increases by 6–13%, confirming broad improvements in originality and fluency (Table 9(Zhang et al., 2020b; Sellam et al., 2020b).
For example, a ZH riddle—“千言万语藏心怀” (lit. “A thousand words hide in the heart”)—exemplifies the character “信” through orthographic metaphor and poetic condensation (Table 20, Row 2), echoing classical radical-based strategies (Tan et al., 2016; Wei and Lee, 2021). This trend represents similar stylistic augmentations across languages, as AOF reduces redundancy (e.g., Self‑BLEU down 40.4% in EN and 42.4% in FR) while increasing diversity (e.g., Distinct‑2 up to 13.5% in EN and 10.6% in ZH).
These cross‑linguistic patterns, quantified in Tables 4 and 9, suggest that AOF enables culture‑attached, cognition‑challenging riddles with higher metaphorical condensation and interpretability. For complete evaluations, see Section A.
@@ 7 AOF Fine-Tuned Evaluations
We visualize cross-language gains in Figure 5 and show their alignment with human judgments in Figure 5, both in Appendix D Fine‑tuning with AOF consistently enhances riddle generation across all five languages by reducing repetition, increasing lexical diversity, and producing more structurally cohesive metaphors. Across the board, RiddleScore increases reflect these quality gains: AR (+57.1%), ZH (+48.3%), EN (+43.4%), FR (+33.7%) and JA (+29.5%) (Table 4). These improvements coincide with major reductions in Self‑BLEU—up to 63.4% for JA and 43.2% for FR—indicating lower reliance on template reuse. Distinct‑2 further supports richer lexical expression, with AR (+18.8%), JA (+31.3%) and FR (+13.3%) seeing the most progress (Table 9). Human evaluation scores for AOF also improved substantially after fine‑tuning (Tables 2 and 2). For example, ZH rose by +15.1%, EN by +16.9%, and JA by +2.1%. FR decreased slightly (−1.3%), while AR maintained its high human evaluation score (4.92). These percentage changes strongly parallel the RiddleScore gains (e.g., ZH: 0.453 $\rightarrow$ 0.728), reinforcing the metric’s validity as a proxy for human perception of creativity, fluency, and cultural fit.
While all languages benefit, fine‑tuning yields especially high returns in languages with deep poetic or idiomatic traditions. For example, in ZH, AOF‑finetuned models generate riddles like “千言万语藏心怀” (“A thousand words hidden in the heart”), whose solution—“信” (message/trust)—demonstrates metaphorical compression grounded in radical‑based inference (Table 20, Row 2). This level of orthographic subtlety is absent in pretrained outputs, underscoring AOF’s value in enabling culturally resonant riddle design.
Methods of prompting vary in consistency: Few-Shot and AOF consistently increase RiddleScore, but Chain-ofThought is inconsistent: significant increases for EN (+48.5%) but negligible for AR (+3. 6%) and JA (0. 0%) - indicating limited generalizability between languages. Only AOF consistently improves human‑aligned and automatic metrics for all languages. Full language‑specific results and examples appear in Section D and Appendices GJ.
9 Fine-Tuned AOF Riddle Comparison to Real World
Across all five languages, fine‑tuned AOF riddles diverge meaningfully from real‑world counterparts by trading formulaic structure for richer metaphor, lexical inventiveness, and cultural depth. Traditional riddles often rely on binary opposites, rhymes, or phonological puns (Gentner, 1983; An, 2023), whereas AOF generations favor conceptual blending (Fauconnier and Turner, 2002), indirect metaphor (Lakoff and Johnson, 1980), and cross‑domain abstraction (Tan et al., 2016).
EN and FR AOF riddles employ echo, shadow, or depth metaphors, including rhythmic phrasings that support recall and poeticity (Encyclopædia Britannica, 2025). For instance, the EN riddle of Table 13, Row 1—“I mirror your thoughts, but never speak”—explores selfhood through contrastive metaphor, absent in real‑world riddles that prefer rhyming antonyms like “shadow/light.” FR AOF riddles follow suit, abandoning “Qu’est‑ce qui” templates for ellipsis‑like phrasing.
In ZH and JA, AOF outputs evoke script‑specific strategies like radical‑based inference and spatial contradiction. The ZH riddle “千言万语藏心怀” (Row 2) reveals “信” (message/trust) through poetic indirection, while the JA riddle “屋根にはいるのに、家にいないものは何?” juxtaposes kanji structure and conceptual space (Sun, 2006; An, 2023).
In AR, fine‑tuned riddles pivot from root‑based puns to symbolic layering, favoring poetic contrasts over mechanical symmetry. As shown in Figure 6, Row 1, metaphors like “a wind that enters but is never welcomed” evoke hospitality norms and classical desert imagery (Al-Khatib, 1988; Antar, 2023; Liu et al., 2022b). For full comparisons and linguistic analysis, see Appendix C.
@@ 10 Conclusion
This paper introduces adaptive originality filtering (AOF), a re-feedback method for improving multilingual riddle generation, pushing models towards semantically new, structurally well-formed, yet culturally embedded, outputs. For five typologically distinct languages, AOF systematically improves human-aligned quality measured by RiddleScore for all five confirming the approach’s universal applicability, regardless of script, form, or model design. These advantages are a byproduct of AOF’s design: AOF discourages revisioning of templates, discourages egocentral phrasing, and trends toward metaphoric, interpretative styles typical for every language’s rhetorical styles. Optimized variants of AOF, besides being better than pretrained generations, by and large are comparable to real-world puzzles by metaphoric richness, especially in very oralistically and visualistically inclined languages Arabic and Chinese. Additionally, AOF generalizes across LLMs, from DeepSeek R1 to GPT-4o and from LLaMA 3.1, in manifesting strong performance across a diversity of generation styles, as well as pretraining corpora. Apart from riddles, this work also suggests that prompting strategies with rejection-based filtering can guide LLMs towards culturally and cognitively compatible results, especially for compositional and figurative tasks.
@@
We limit our experiment to the BiRdQA corpus, comprised of 6,614 English and 8,751 Chinese multi-choice riddles. Though genre-various, its figurative concentration limits generalizability to larger creative tasks (e.g., allegory or storytelling). Our five-lingual evaluation extends over EN–ZH–AR–JA–FR, but omits lower-resource or more-morphologically challenging languages like Finnish or Swahili.
@@
We uniformly set decoding hyperparameters (e.g., temperature, number of tokens) to allow for comparison, but possibly suppress interactions between prompts and parameters. Filtering by MiniLM in AOF targets semantic novelty, but cosine similarity may overlook certain subtle redundancies, especially where languages are morphologically diverse or idiomatic.
@@
Our GPT‑4o fine-tuning uses BiRdQA’s multiple-choice setup, boosting structural fluency but potentially biasing toward riddles that privilege explicit clarity over conscious ambiguity. While stylistic refinement shows up by metrics such as Self‑BLEU, Distinct‑2 and RiddleScore, more detailed downstream measurements such as solver accuracy and difficulty calibration are left to future research.
@@
Human judgments were made by native or proficient speakers from five languages employing standard rubrics. This guarantees cultural anchoring but sample size and analysis by inter‑annotator agreement were restricted by resources. To evaluate creativity, fluency, and cultural fit, RiddleScore, tested against these ratings, yields an interpretable proxy, albeit a proxy that doesn’t register longer‑term aspects like memorability, interest, or difficulty to solve.
@@
This research assesses riddle-making within five languages, including English, Chinese, Japanese, Arabic, and French, selected to be typologically diverse and with resources to draw from. Although this gives a wide cultural span, the dataset and prompts come from internet-based corpora and so might not capture perfectly idiomatic richness from less represented populations. Certain metaphorical or rhetorical patterns might be overly represented within English or less developed within other languages even with our balancing qualitative with quantitative assessment.
@@
Procedurally generated riddles may resemble publicly known riddles from folk sources or online corpora. As described in Sections 3–4, Adaptive Originality Filtering (AOF) mitigates this risk by rejecting outputs with high semantic similarity to reference data. Nonetheless, we caution against deploying outputs in commercial settings without additional originality verification. AI assistants (e.g., ChatGPT) were also used to support code development and manuscript preparation. During implementation, LLMs aided in debugging and optimizing evaluation scripts (e.g., for RiddleScore and Distinct-2). In writing, AI was used for linguistic refinement, including phrasing, transitions, and caption clarity. All methodological contributions, analysis, and final revisions were conducted by the authors.
@@
These data have no personally identifiable information (PII). The riddles are anonymized and cast as general-knowledge metaphors. The fine-tuning followed OpenAI’s API regulations, token constraints, and safety limits, and never involved user-submitted or private material.
Human Evaluation and Metric Ethics
Human ratings were made by native or expert speakers with standardized rubrics, allowing for culturally sensitive evaluations. Model IDs were blinded to help decrease bias. RiddleScore, tested against these human ratings, provides a formalized proxy to creativity, fluency, and cultural fit but doesn’t assess engagement, memorability, or difficulty for solvers.
@@
While generation of riddles is a low-risk task, their creative uncertainty might be exploited to spread misinformation or to manipulate culturally sensitive information. We advise against using them in high-stakes educational, psychological, or legal applications without interpretability controls and human review.
Appendix A Appendix: AOF Fine-tuned language Evaluations